If 20 ml of 0.10m naoh was used for the titration of aspirin, calculate the mass of the aspirin.


1
Expert's answer
2019-10-31T07:29:20-0400

Answer on Question #97578 – Chemistry – Other


Task:

If 20 ml of 0.10M NaOH was used for the titration of aspirin, calculate the mass of the aspirin.


Solution:

Aspirin = C9H8O4 = Asp

1) Balance equation for the titration of aspirin:

C9H8O4 + 2NaOH = C7H5O3Na + C2H3O2Na + H2O.

For simplicity, let's rewrite the equation as:

Asp + 2NaOH = NaSal + NaAc + H2O.

2) By equation:

n(Asp) = n(NaOH) / 2.

3) Calculate the moles of NaOH used in the reaction:

n(NaOH) = C(NaOH)*V(NaOH) = 0.10M * 0.02L = 0.002 mol.

4) Calculate the moles of aspirine:

n(Asp) = n(NaOH) / 2 = 0.002 mol / 2 = 0.001 mol.

5) Calculate the mass of aspirin:

m(Asp) = n(Asp)*M(Asp);

M(Asp) = M(C9H8O4) = 9*Ar(C) + 8*Ar(H) + 4*Ar(O) = 9*12.0096 + 8*1.0078 + 4*15.999 = 180.1448 (g/mol);

m(Asp) = 0.001 mol * 180.1448 g/mol = 0.18 g.

m(Asp) = 0.18 g.


Answer: m(aspirin) = 0.18 g.